# Mortality Modelling using Generalized Estimating Equations

### 20 Jan 2024 v1

@@ 2.3 Single Population Analysis - Czech Republic (Males Age 20-80)
Focusing on Czech males aged 20 to 80 during 1991-2010, we employ GEE model 3 with independence (m1), exchangeable (m2), AR(1) (m3), and unstructured (m4) correlation structures. Table 1 shows QIC values for model selection. Despite close QICs, m1 and m2 feature lower Corrected Information Criterion (CIC) values. We choose m2 as the final model for predicting mortality rates in Czech males from 2011 to 2019 (pre-COVID-19).
@@
m1 (Independence) $5.01$ $493$ $-0.251$ $2.27\times 10^{-23}$ $-645$
m2 (Exchangeable) $5.01$ $493$ $-0.251$ $2.46\times 10^{-23}$ $-647$
m3 (AR(1)) $5.01$ $493$ $-0.251$ $2.01\times 10^{-14}$ $-647$
m4 (Unstructured) $5.09$ $493$ $-0.255$ $2.87\times 10^{-5}$ $-1000$
@@ 2.4 Multi-Population Analysis - Austria and Czech Republic (Females and Males, Age 20-80)
This section explores mortality rates for Austria and Czech Republic females and males aged 20 to 80 during 1991-2010, serving as the training set. Utilizing three GEE models (m1, m2, m3) with independence, exchangeable, and AR(1) correlation structures, Table 2 displays QIC values for model selection. Opting for m2, incorporating an exchangeable structure, this final model predicts mortality rates for the four populations during 2011-2019. While QIC and CIC values for Independence and Exchangeable structures are comparable, the exchangeable structure is deemed more suitable for ASDRs due to its assumption that the correlation between data points remains consistent over time, aligning with the nature of ASDR trends.
@@ 3 Conclusion
This study has explored the utility of GEEs as an efficient tool for modelling mortality rates. By conducting a thorough analysis of longitudinal mortality datasets, GEEs have demonstrated their computational simplicity, particularly in effectively handling categorical data and correlated observations. However, we need to be careful about the inherent limitations in GEE modelling. The absence of a fully specified likelihood function and constraints in likelihood-based methods may impact the extent of statistical analysis achievable with GEEs. Additionally, cautious interpretation is warranted, especially concerning empirical-based standard errors, which may lead to underestimation, particularly in smaller sample sizes. Striking a balance between computational efficiency and nuanced model specifications becomes pivotal when utilizing GEEs in mortality rate modelling.

### 23 Jan 2024 v2

@@ 2.3 Single Population Analysis - Czech Republic (Males Age 20-80)
Focusing on Czech males aged 20 to 80 during 1991-2010, we employ GEE model 3 with independence (geeInd), exchangeable (geeEx), AR(1) (geeAr1), and unstructured (geeUns) correlation structures. Table 1 shows QIC values for model selection. Despite close QICs, geeInd and geeEx feature lower Corrected Information Criterion (CIC) values. We choose geeEx as the final model for predicting mortality rates in Czech males from 2011 to 2019 (pre-COVID-19).
@@
geeInd $5.01$ $493$ $-0.251$ $2.27\times 10^{-23}$ $-645$
geeEx $5.01$ $493$ $-0.251$ $2.46\times 10^{-23}$ $-647$
geeAr1 $5.01$ $493$ $-0.251$ $2.01\times 10^{-14}$ $-647$
geeUns $5.09$ $493$ $-0.255$ $2.87\times 10^{-5}$ $-1000$
@@ 2.4 Multi-Population Analysis - Austria and Czech Republic (Females and Males, Age 20-80)
This section explores mortality rates for Austria and Czech Republic females and males aged 20 to 80 during 1991-2010, serving as the training set. Utilizing three GEE models (geeInd, geeEx, geeAr1) with independence, exchangeable, and AR(1) correlation structures, Table 2 displays QIC values for model selection. Opting for geeEx, incorporating an exchangeable structure, this final model predicts mortality rates for the four populations during 2011-2019. While QIC and CIC values for Independence and Exchangeable structures are comparable, the exchangeable structure is deemed more suitable for ASDRs due to its assumption that the correlation between data points remains consistent over time, aligning with the nature of ASDR trends.
@@ 3 Conclusion
This study explores the utility of GEEs as an efficient tool for modelling mortality rates. By conducting a thorough analysis of longitudinal mortality datasets, GEEs demonstrate their computational simplicity, particularly in effectively handling categorical data and correlated observations. However, we need to be careful about the inherent limitations in GEE modelling. The absence of a fully specified likelihood function and constraints in likelihood-based methods may impact the extent of statistical analysis achievable with GEEs. Additionally, cautious interpretation is warranted, especially concerning empirical-based standard errors, which may lead to underestimation, particularly in smaller sample sizes. Striking a balance between computational efficiency and nuanced model specifications becomes pivotal when utilizing GEEs in mortality rate modelling.